• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 08:00:04 on January12,2002

    Entry number 80134

    keyword=Shift summary

    00:00 Beginning of shift, MCC still working on various problems and beam tune. ROC4 needs to be reset during the next access.
    00:36 MCC called. They will send us pulsed beam to trouble shoot the problem.
    00:47 MCC called. They will send us CW beam to see beam on target.
    01:00 MCC called to inform us that the value on the radiator was reading zero. Since the slow target controls are not working, we cannot confirm the status of the radiator.
    01:10 MCC called. They wanted permission to mask the target FSD. I paged Arun Saha since they need his promission.
    01:55 Doug talked to MCC, and they decided to try to send beam straight to the hall without going through the compton Chicane.
    02:21 MCC is sending pulsed beam.
    02:44 MCC will attempt to deliver CW beam.
    03:00 MCC reports that they can now deliver CW beam, but there are still some problems with BPMs upstream of the target.
    03:50 Doug increased the beam current to 2.86 microamps to see if we can see the beam on target.
    04:10 Doug performed a Harp scan 1455. Scan was good. See Halog # 80124. However Fast Feedback is not working, and an energy measurement will not be performed during this shift.
    04:40 Asked for a controlled access to reset the VME crates and install a filter on S0.
    05:30 Access to Hall A
    05:40 VME crate reset and filter installed
    06:10 Work completed in the hall, but the fastbus crate needs to be recycled.
    06:43 Access to Hall to recycle the fastbus crate.
    07:05 L_BTM VDC HV tripped to zero. We reset the VDC crates, and the VDC level is back to 4.0 kV. Note: The VDC HV rarely trips.
    8:00 At the end of shift, the accelerator still needs to fix the fastfeed back, and the fastbus (ROC2) still needs to be fixed.